# Particle Spectra in the Integrated Hydrokinetic Model at RHIC BES Energies

### v1

@@ II.0.3 Hydrodynamic expansion, particlization, and hadron cascade
After thermalization is achieved, iHKM follows a classical path of hybrid models. At $\tau_{th}$, the system expands via viscous hydrodynamic equations, which we solve numerically using the vHLLE code [28]. Once the system becomes dilute and reaches the critical energy density $\epsilon_{\text{sw}}$, the evolution switches back to the UrQMD cascade. To determine the particlization hypersurface of constant energy density, we employ the Cornelius routine [29, 30]. The particlization of liquid into hadron resonant gas is realized via a Cooper-Frye-like algorithm
$N^{i}=\int d^{3}p\int\frac{d\sigma_{\mu}p^{\mu}}{p_{0}}f_{i}(x,p).$ (14)
Here $N^{i}$ is a number of hadrons of species $i$ emitted from the element of hypersurface $d\sigma_{\mu}$ and $f_{i,eq}(x,p)$ is a distribution function. For high-energy collisions, the Bose-Einstein or Fermi-Dirac equilibrium distribution function with a small correction is a fairly good assumption. However, if the equilibrium is not reached, which we expect at several GeV experiments, two components must be considered. Following Eqs. (11), (12) in relaxation time approximation we obtain
@@
where $f_{\texttt{eq.}}(x,p)$ is a near-equilibrium component constructed from the output of the hydrodynamic component, and $f_{\texttt{n.eq.}}(x,p)$ is a far from equilibrium distribution function constructed from UrQMD with a kernel (5).
@@ III.0.1 Free parameters
To evaluate the sensitivity to the softening of the equation of state, we consider two qualitatively distinct models: the chiral EoS (CO) [22], which features a cross-over transition between the quark-gluon plasma and the hadron resonance gas, and an EoS with a first-order phase transition (PT) [23]. For each EoS, we perform simulations varying the free parameters as follows:
  • $\tau_{0}$ $\left[\texttt{fm}/c\right]$ start of the thermalization stage
  • $\eta/s$ $\left[1\right]$ shear viscosity to entropy density ratio. Scaling factor in shear stress tensor $\pi^{\mu\nu}$
  • $\tau_{\text{th}}$ $\left[\texttt{fm}/c\right]$ end of the thermalization stage
@@
  • $\tau_{\text{rel}}$ $\left[\texttt{fm}/c\right]$ relaxation time
  • $R$ $\left[\texttt{fm}\right]$ gaussian smearing parameter in kernel (5)
To simplify the calibration procedure and reduce the number of free parameters, we fix the relaxation time to its maximum permitted value, effectively transferring all remaining flexibility to the parameter $\tau_{\text{th}}$.
$\tau_{\text{rel}}=\tau_{\text{th}}-\tau_{0}.$ (16)
The Gaussian smearing parameter is fixed at a reasonable, though arbitrary, value of $R=0.5$ fm for all initial conditions, as there is no clear physical justification for varying it between different experimental setups. Consequently, the smoothness of the initial state in our model is governed entirely by the thermalization stage. In summary, only four parameters, along with the choice of equation of state, are treated as free.
III.0.2 Sensitivity of spectra to the parameters
To explore the correlations between model parameters and observables, we generate 750 simulation sets for Au-Au collisions at $\sqrt{s_{NN}}=14.5$ GeV within the $20-30\%$ centrality class. The parameter values for each simulation are sampled using Latin hypercube sampling within the ranges specified in Table 1. In this analysis, we allow variations in both the $R$ and $\tau_{\text{rel}}$ parameters, while ensuring that the constraint given by Eq. (16) for $\tau_{th}$.
@@ III.0.2 Sensitivity of spectra to the parameters
Table 1: Minimal and maximal allowed values of model parameters in Latin hypercube.
For each parameter set, we generate 40 initial conditions using UrQMD. To improve statistical accuracy, we simulate 200 afterburner events on the particlization hypersurface for each initial condition, resulting in a total of 8,000 events per parameter set. We then construct transverse momentum spectra for pions, kaons, and (anti)protons. Fig. 1 shows the transverse momentum spectra at midrapidity, averaged over 750 parameter sets, with a one-standard-deviation range for both equations of state. On average, the midrapidity particle yields are higher in calculations with the softer phase-transition equation of state. However, this scenario also leads to weaker transverse flow, which is reflected in the high-$p_{T}$ spectra. It is important to note that the mean values of the parameters listed in Table 1 do not correspond to the best-fit values for the experimental data. As a result, the spectra corresponding to the mean parameter set may lie outside the one-standard-deviation ranges shown (blue and red shaded bands) in Fig. 1.
Figure 1: $p_{T}$ spectra for $\pi^{-}$, $p$, and $\bar{p}$ (from left to right) averaged over $750$ simulations, with one standard deviation from the mean values, using sets with random parameters within the ranges presented in Table 1. Data correspond to Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV and 20-30% centrality class.
@@
The sign of the correlation matrix element indicates whether the parameter and observable are correlated ($+$) or anticorrelated [^3] An increase in the parameter value leads to a decrease in the observable value. (--), while the absolute value reflects the strength of the correlation, with “0” representing no correlation and “$\pm$1” corresponding to a strong (anti) correlation.
@@
Figure 2: Correlation coefficients between the five model parameters and the transverse momentum spectra at low, intermediate, and high $p_{\text{T}}$ for $\pi^{-}$, $p$, and $\bar{p}$ are presented. The results for $\pi$ mesons and protons (shown in the two tables on the left) are very similar for both equations of state; therefore, only the crossover case is shown. For antiprotons (shown in the two plots on the right), we present the results for both scenarios.
@@
We construct correlation matrices for six particle species: $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$, using both equations of state. However, since the results for pions and kaons are very similar, Figure 2 presents only the correlations for negatively charged pions, protons, and antiprotons. Additionally, we show correlations only for antiprotons for the soft EoS with a first-order phase transition, as the results for all other cases are qualitatively and quantitatively similar to those of the crossover scenario.
@@
One key conclusion from Figure 2 is that the common smoothing procedure for initial conditions (encoded in the parameter $R$) plays a crucial role in hydrodynamic simulations, which limits our ability to infer properties of strongly interacting matter in heavy-ion collisions. A second observation is that the parameter $\tau_{\text{rel}}$ behaves similarly to $\tau_{0}$, although its effect on the observables is less pronounced. We also find that, unlike other hadron species, antiproton yields are particularly sensitive to the switching energy density $\epsilon_{\text{sw}}$, which marks the transition from hydrodynamics to transport. This sensitivity is also clearly demonstrated in [31]. It may be explained by the suppression of (anti)baryon yields in transport models like UrQMD [32, 33] or SMASH [34] used as afterburner on the final stage of the system’s evolution. This issue has been addressed in numerous papers to which we refer the reader [35, 36, 37].
@@
When comparing the correlation matrices for antiprotons, we observe that in the case of a softer equation of state (PT EoS), the parameters defining the pre-thermal evolution (i.e., $\tau_{0}$, $\tau_{\text{rel}}$, $R$) play a more significant role than in the case of the crossover equation (CO EoS). This is reflected in the larger absolute values of the corresponding correlation coefficients. In contrast, the particlization energy density ($\epsilon_{\text{sw}}$), which defines the late-stage evolution, becomes more influential when a stiffer equation of state is employed.
@@ III.0.3 Model finetuning
In this study, we use particle spectra to calibrate the model’s free parameters. As shown in Fig. 2, the initiation of the thermalization stage, denoted by $\tau_{0}$, emerges as one of the most critical parameters. It is anticipated that in midrapidity, thermalization can begin even before the nuclei fully overlap, at $\tau_{\text{overlap}}$ (see Eq. (1)). Our numerical analysis indicates that the best agreement with experimental pion spectra is achieved when $\tau_{0}$ is approximately $0.75\tau_{\text{overlap}}$. The conclusion of the thermalization phase has a significantly weaker influence on observables and, therefore, is determined with greater uncertainty. Nonetheless, it generally occurs at or slightly after the nuclei’ overlapping time.
@@
Table 2: Parameters of iHKM providing the best description of bulk observables for Au+Au collisions at $\sqrt{s_{NN}}$ = 7.7, 11.5, 14.5, 19.6, 27.0, 39.0 GeV.
$\sqrt{s_{NN}}$ EoS R $\tau_{0}$ $\tau_{th}$ $\eta$ / s $\epsilon_{\text{sw}}$
7.7 PT 0.5 2.7 3.3 0.08 0.35
7.7 CO 0.5 2.5 3.3 0.08 0.50
11.5 PT 0.5 2.0 2.6 0.08 0.35
11.5 CO 0.5 1.8 2.6 0.08 0.50
14.5 PT 0.5 1.4 2.0 0.08 0.35
14.5 CO 0.5 1.3 2.3 0.08 0.50
19.6 PT 0.5 1.0 1.6 0.08 0.35
19.6 CO 0.5 1.0 1.6 0.08 0.50
27.0 PT 0.5 0.8 1.4 0.08 0.35
27.0 CO 0.5 0.8 1.4 0.08 0.50
39.0 PT 0.5 0.6 1.2 0.08 0.35
39.0 CO 0.5 0.6 1.6 0.08 0.50
Since low-$p_{T}$ spectra show a weak dependence on $\eta/s$, we calibrate the viscosity by focusing on relatively high-$p_{T}$ spectra (around $1$ GeV/$c$) after fixing the thermalization stage parameters. Furthermore, we account for differences in the scale and slope of particle spectra when using different equations of state (see Fig.2). Because the antiproton spectra are particularly sensitive to $\eta/s$ and $\epsilon_{sw}$, we use experimental data for antiprotons to estimate these parameters. Although their values differ for the two equations of state, we keep them the same while varying the collision energy $\sqrt{s_{NN}}$.
@@ III.0.4 Momentum spectra
In this section, we present the iHKM results for transverse momentum spectra in Au+Au collisions at energies of 7.7, 11.5, 14.5, 19.6, 27, and 39 GeV, across two centrality classes: 20-30% and 0-5%, as shown in Figs. 4 and 5. The final model parameters are summarized in Table 2. The $\pi^{-}$ spectra for the 20-30% centrality class were used to calibrate the model parameters. To improve the proton-to-antiproton ratio, the $\epsilon_{sw}$ parameter was allowed to change between different equations of state, but it was kept constant across collision energies. We also do not change model parameters for different centrality classes.
It is important to note that at RHIC BES, centrality classes are defined by the multiplicity of charged particles at midrapidity. In hydrodynamic simulations, using minimal-bias events for model calibration is inefficient. Therefore, we rely on the multiplicities of UrQMD events used for the initial conditions, extending them to longer times of up to 200 fm/$c$.
In Fig. 3, we demonstrate how our inferred value of $\tau_{0}$, derived from pion spectra, depends on the collision energy $\sqrt{s_{NN}}$. We found that it lies within a range of ±$\pm 0.2$ fm/$c$ from the value of $\tau_{\text{overlap}}$, scaled by a factor of 0.75. It is important to note that assuming thermalization begins at a fixed proper time $\tau$ is a simplification that may only be valid in the central region of the system. Upon analyzing the spectra presented in Figs. 5 and 4, several key trends emerge. First, the difference between the two equations of state is relatively small and becomes negligible at higher collision energies. This behavior is also reflected in the model parameters listed in Table 2.
@@ IV Summary
In this paper, we employ a recently developed, extended version of the iHKM model to describe the transverse momentum spectra of light hadrons in the Beam Energy Scan (BES) program at the Relativistic Heavy Ion Collider (RHIC). Our primary focus is to explore the relationship between the model’s free parameters and experimental observables. In particular, we investigate how variations in the thermalization time scales and the switching conditions between the hydrodynamic and hadronic stages influence particle spectra and transverse momentum distributions. Simulations are performed using the two distinct equations of state (EoS): one featuring a crossover transition in baryon-rich matter between quark-gluon plasma and hadron gas, and the other incorporating a first-order phase transition.
@@
We also find that both equations of state yield similarly accurate descriptions of the transverse momentum spectra when adjustments are made to the relaxation time scale. At higher collision energies within the BES range, differences between the results for spectra in the two scenarios become negligible. The analysis of elliptic flow is necessary.
The most pronounced difference between the two EoS scenarios emerges at $\sqrt{s_{NN}}=7.7$ GeV, particularly in the yields of protons and kaons. This discrepancy is likely due to the differing trajectories the systems follow on the $T$$\mu_{B}$ phase diagram during their evolution. The existence of a phase transition line, along with the choice of switching energy density $\epsilon_{\text{sw}}$ to the hadronic afterburner, plays a critical role in determining the parameters at chemical freeze-out. A more comprehensive analysis of additional bulk observables, especially of elliptic flow, within the iHKM framework is planned for future work.

### v2

@@ II.0.3 Hydrodynamic expansion, particlization, and hadron cascade
After thermalization is achieved, iHKM follows a classical path of hybrid models. At $\tau_{th}$, the system expands via viscous hydrodynamic equations, which we solve numerically using the vHLLE code [49]. Once the system becomes dilute and reaches the critical energy density $\epsilon_{\text{sw}}$  [^4] As in our previous work [31], we use the energy density obtained from the hydrodynamic tensor $T^{\mu\nu}_{\rm hydro}$ according to Eq. (8)., the evolution switches back to the UrQMD cascade. To determine the particlization hypersurface of constant energy density, we employ the Cornelius routine [50, 51]. The particlization of liquid into hadron resonant gas is realized via a Cooper-Frye-like algorithm
$N_{i}=\int d^{3}p\int\frac{d\sigma_{\mu}p^{\mu}}{p_{0}}f_{i}(x,p).$ (13)
Here $N^{i}$ is a number of hadrons of species $i$ emitted from the element of hypersurface $d\sigma_{\mu}$ and $f_{i}(x,p)$ is a distribution function. For high-energy collisions, the Bose-Einstein or Fermi-Dirac equilibrium distribution function with a small correction is a fairly good assumption. However, if the equilibrium is not reached, which we expect at several GeV experiments, two components must be considered. Following Eqs. (10), (11) in relaxation time approximation we obtain
@@
where $f_{\texttt{eq.}}(x,p)$ is a near-equilibrium component constructed from the output of the hydrodynamic component, and $f_{\texttt{n.eq.}}(x,p)$ is a far from equilibrium distribution function constructed from UrQMD with a kernel (5). Consequently, we sampled hadrons for this component from the same UrQMD event that was used for initial dynamics.
@@ III.0.1 Free parameters
To assess the sensitivity to the softening of the equation of state, we consider two qualitatively distinct models: the chiral EoS [38], which exhibits a crossover (CO) transition between the quark-gluon plasma and the hadron resonance gas, and an EoS featuring a first-order phase transition (PT) [39]. For each EoS, we conduct simulations while systematically varying the following free parameters introduced in the previous section:
  • $\tau_{0}$ $\left[\text{fm}/c\right]$ - start of the thermalization stage
  • $\eta/s$ $\left[1\right]$ - shear viscosity to entropy density ratio. Scaling factor in shear stress tensor $\pi^{\mu\nu}$
  • $\tau_{\text{th}}$ $\left[\text{fm}/c\right]$ - end of thermalization
  • $\epsilon_{\text{sw}}$ $\left[\text{GeV/fm}^{3}\right]$ - hydrodynamic energy density corresponding to the transition to the hadron afterburner stage
  • $\tau_{\text{rel}}$ $\left[\text{fm}/c\right]$ - relaxation time
  • $R$ $\left[\text{fm}\right]$ - gaussian smearing parameter in kernel (5)
To simplify the calibration procedure and reduce the number of free parameters, we fix the relaxation time to its maximum permitted value, effectively transferring all remaining flexibility to the $\tau_{\text{th}}$ parameter.
$\tau_{\text{rel}}=\tau_{\text{th}}-\tau_{0}.$ (15)
III.0.2 Sensitivity of spectra to the parameters
@@
[…]
@@
Table 1: Ranges of model parameters used in the Latin hypercube sampling.
[…]
Figure 1: Transverse momentum ($p_{T}$) spectra for $\pi^{-}$, $p$, and $\bar{p}$ (from left to right) averaged over $750$ simulations, with one standard deviation from the mean values, using sets with random parameters within the ranges presented in Table 1. Data correspond to Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV and 20-30% centrality class.
@@
The sign of the correlation matrix element indicates whether the parameter and observable are correlated (++) or anticorrelated [^5] An increase in the parameter value leads to a decrease in the observable value. (-), while the absolute value reflects the strength of the correlation, with “0” representing no correlation and $\pm$1” corresponding to a strong (anti) correlation.
@@
Figure 2: Correlation coefficients between the five model parameters and the transverse momentum spectra at low, intermediate, and high $p_{\text{T}}$ for $\pi^{-}$, $p$, and $\bar{p}$ in $20-30\%$ centrality Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV. The results for $\pi^{-}$ mesons and protons (shown in the two tables on the left) are very similar for both equations of state; therefore, only the crossover case is shown. For kaons, the correlation matrices are similar to those of pions. For antiprotons (shown in the two plots on the right), we present the results for both scenarios.
@@
We investigate correlation matrices for six particle species: $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$, using both equations of state. However, since the results for pions and kaons are very similar, in Figure 2 we present only the correlations for negatively charged pions, protons, and antiprotons. Additionally, we show correlations only for antiprotons for the soft EoS with a first-order phase transition, as the results for all other cases are qualitatively and quantitatively similar to those of the crossover scenario.
@@
One key conclusion from Fig. 2 is that the commonly used smoothing procedure for initial conditions, encoded in the parameter $R$, strongly influences hydrodynamic simulations and limits our ability to extract properties of the matter created in heavy-ion collisions. A second observation is that the parameter $\tau_{\rm rel}$ behaves similarly to $\tau_{0}$, although its impact on the observables is less pronounced. We also find that, unlike other hadron species, antiproton yields are particularly sensitive to the switching energy density $\epsilon_{\text{sw}}$, which marks the transition from hydrodynamics to transport. This sensitivity, also demonstrated in Ref. [55], can be attributed to the suppression of (anti)baryon yields in transport models such as UrQMD [56, 32] or SMASH [57], which are employed as afterburners in the final stage of the system’s evolution. This issue has been addressed in numerous studies, to which we refer the reader [58, 59, 60].
@@
When comparing the correlation matrices for antiprotons, we observe that in the case of a softer equation of state (PT EoS), the parameters defining the pre-thermal evolution (i.e., $\tau_{0}$, $\tau_{\text{rel}}$, $R$) play a more significant role than in the case of the crossover equation (CO EoS). This is reflected in the larger absolute values of the corresponding correlation coefficients. In contrast, the particlization energy density ($\epsilon_{\text{sw}}$), which defines the late-stage evolution, becomes more influential when a stiffer equation of state is employed. Finally, we note that Pearson correlations capture only linear relationships. The results may depend on collision energy, particularly at low energies where the system spends longer in the transition region of the QCD phase diagram and the differences between equations of state are more pronounced.
@@ III.0.3 Model finetuning
Since low-$p_{T}$ spectra exhibit only a weak dependence on $\eta/s$ [61], we fix this parameter at a typical value of $\eta/s=0.08$ [62], deferring its detailed tuning to future studies with a more comprehensive analysis. Because antiproton spectra are particularly sensitive to the switching energy density $\epsilon_{\rm sw}$, we use experimental antiproton data to constrain this parameter. Although the optimal values of $\epsilon_{\rm sw}$ differ for the two equations of state, we do not vary them when changing the collision energy $\sqrt{s_{NN}}$.
The smoothing parameter $R$ appears in various forms across most dynamical models with event-by-event fluctuating initial conditions, yet its value remains poorly constrained. For example, Ref. [35] uses R=0.5fmR=0.5\ $\text${fm}, Ref. [63] adopts 1fm1\ $\text${fm}, and Ref. [64] treats it as a free parameter. In iHKM, smoothing is further influenced by the thermalization stage, and there is no strong evidence to vary $R$ across different collision energies. Therefore, we set $R$ to the smallest reasonable value of 0.5fm0.5\ $\text${fm}. Smaller values result in overly spiky initial distributions, which can destabilize the early hydrodynamic evolution.
The final iHKM parameters for different RHIC BES energies are summarized in Table 2. In Fig. 3, we show how the inferred value of $\tau_{0}$, obtained from pion spectra, depends on the collision energy $\sqrt{s_{NN}}$. We find that $\tau_{0}$ lies within $\pm 0.2$ fm/$c$ of the scaled overlap time, $0.75\,\tau_{\text{overlap}}$. This indicates that, in our model, thermalization begins slightly before the nuclei fully overlap, but still concludes slightly afterward. It is important to note that assuming thermalization starts at a fixed proper time $\tau$ is a simplification, which may be valid only in the central region of the system.
@@
Table 2: Optimized iHKM parameters that provide the best description of bulk observables in Au+Au collisions at $\sqrt{s_{NN}}=7.7,11.5,14.5,19.6,27.0$, and $39.0$ GeV. The remaining parameters are fixed at $R=0.5$ fm and $\eta/s=0.08$.
$\sqrt{s_{NN}}$ (GeV) EoS $\tau_{0}$ (fm/$c$) $\tau_{th}$ (fm/$c$) $\epsilon_{\text{sw}}$ (GeV/fm3)
7.7 PT 2.7 3.3 0.35
7.7 CO 2.5 3.3 0.50
11.5 PT 2.0 2.6 0.35
11.5 CO 1.8 2.6 0.50
14.5 PT 1.4 2.0 0.35
14.5 CO 1.3 2.3 0.50
19.6 PT 1.0 1.6 0.35
19.6 CO 1.0 1.6 0.50
27.0 PT 0.8 1.4 0.35
27.0 CO 0.8 1.4 0.50
39.0 PT 0.6 1.2 0.35
39.0 CO 0.6 1.6 0.50
@@ III.0.4 Momentum spectra
Now, we present the iHKM results for transverse momentum spectra in Au+Au collisions at energies of 7.7, 11.5, 14.5, 19.6, 27, and 39 GeV, across two centrality classes: 20-30% and 0-5%, as shown in Figs. 4 and 5. Analysis of the spectra shown in Figs. 5 and 4 reveals several key trends. First, the differences between the results for the two equations of state are relatively small and become negligible at higher collision energies. This trend is also reflected in the model parameters listed in Table 2. The observed behavior can be attributed to the similar description of the QGP phase in both equations of state, while the main differences arise near the transition to the hadron resonance gas. At higher collision energies, the rapid expansion of the system makes this transition occur too quickly for the differences to significantly affect the spectra, whereas at lower energies the differences are more pronounced.
@@ IV Summary
In this paper, we employ a recently developed, extended version of the iHKM model to describe the transverse momentum spectra of light hadrons in the Beam Energy Scan program at the Relativistic Heavy Ion Collider. Our primary focus is to explore the relationship between the model’s free parameters and experimental observables. In particular, we investigate how variations in the thermalization time scales and the switching conditions between the hydrodynamic and hadronic stages influence particle spectra and transverse momentum distributions. Simulations are performed using the two distinct equations of state: one featuring a crossover transition in baryon-rich matter between quark-gluon plasma and hadron gas, and the other incorporating a first-order phase transition.
Our analysis reveals that the duration of the thermalization stage is approximately 1 fm/$c$ across all RHIC BES energies in the range $\sqrt{s_{NN}}=7.7$$39$ GeV. However, thermalization begins earlier at lower collision energies. Specifically, our numerical results indicate that thermalization commences slightly before the two colliding nuclei fully overlap, assuming they propagate with their initial rapidities; this corresponds to $\tau_{0}\approx 0.75\,\tau_{\text{overlap}}$ for both considered equations of state.
We also find that both equations of state provide comparably accurate descriptions of the transverse momentum spectra, once the relaxation timescale is appropriately adjusted. At higher collision energies within the BES range, differences between the two scenarios in the spectra become negligible. A detailed analysis of other bulk observables is therefore required.